Cost of Common Equity and WACC

Palencia Paints Corporation has a target capital structure of 45% debt and 55% common equity, with no preferred stock. Its before-tax cost of debt is 13% and its marginal tax rate is 40%. The current stock price is P0 = $33.50. The last dividend was D0 = $2.50, and it is expected to grow at a 7% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.
Answer #1

Cost of common equity=(D1/Current price)+Growth rate

=[(2.5*1.07)/33.5]+0.07

=14.99%(Approx).

Cost of debt aftertax=13*(1-tax)

=13*(1-0.4)=7.8%

WACC=Respective costs*Respective weight

(14.99*0.55)+(0.45*7.8)

=11.75%(Approx).